# DiReCT: Diagnostic Reasoning for Clinical Notes via Large Language Models

### v2

@@ Abstract
Large language models (LLMs) have recently showcased remarkable capabilities, spanning a wide range of tasks and applications, including those in the medical domain. Models like GPT-4 excel in medical question answering but may face challenges in the lack of interpretability when handling complex tasks in real clinical settings. We thus introduce the diagnostic reasoning dataset for clinical notes (DiReCT), aiming at evaluating the reasoning ability and interpretability of LLMs compared to human doctors. It contains 511 clinical notes, each meticulously annotated by physicians, detailing the diagnostic reasoning process from observations in a clinical note to the final diagnosis. Additionally, a diagnostic knowledge graph is provided to offer essential knowledge for reasoning, which may not be covered in the training data of existing LLMs. Evaluations of leading LLMs on DiReCT bring out a significant gap between their reasoning ability and that of human doctors, highlighting the critical need for models that can reason effectively in real-world clinical scenarios Code are available https://github.com/wbw520/DiReCT. Data will be released through PhysioNet..
@@ 1 Introduction
Despite the advancements, interpretability is critical, particularly in medical NLP tasks [Liévin et al., 2024]. Some studies assess this capability over medical QA [Pal et al., 2022, Li et al., 2023, Chen et al., 2024] or natural language inference (NLI) [Jullien et al., 2023]. Putting more attention on interpretability, they use relatively simple tasks as testbeds, taking short text as input. However, tasks in real clinical settings can be more complex [Gao et al., 2023a]. As shown in Figure 1, a typical diagnosis requires comprehending and combining various information, such as health records, physical examinations, and laboratory tests, for further reasoning of possible diseases in a step-by-step manner following the established guidelines. This observation suggests that both perception, or reading, (e.g., finding necessary information in medical record) and reasoning (determining the disease based on the observations) should be counted when evaluating interpretability in LLM-based medical NLP tasks.
@@
To underscore the challenge offered by our dataset, we evaluate a simple AI-agent based baseline [Xi et al., 2023, Tang et al., 2023] that utilizes the knowledge graph to decompose the diagnosis into a sequence of diagnoses from a smaller number of observations. Our experimental findings indicate that current state-of-the-art LLMs still fall short of aligning well with human doctors.
Contribution. DiReCT offers a new challenge in diagnosis from a complex clinical note with explicit knowledge of established guidelines. This challenge aligns with a realistic medical scenario that doctors are experiencing. In the application aspect, the dataset facilitates the development of a model to support doctors in diagnosis, which is error-prone [Middleton et al., 2013, Liu et al., 2022]. From the technical aspect, the dataset can benchmark models’ ability to read long text and find necessary observations for multi-evidence entailment tree reasoning. As shown in Figure 3, this is not trivial because of the variations in writing; superficial matching does not help, and medical knowledge is vital. Meanwhile, reasoning itself is facilitated by the knowledge graph. The model does not necessarily have the knowledge of diagnostic guidelines. With this choice, the knowledge graph explains the reasoning process, which is also beneficial when deploying such a diagnosis assistant system in practical uses.
@@ 3.2 Diagnostic Knowledge Graph
Our knowledge graphs $\mathcal{K}=\{\mathcal{K}_{i}\}_{i}$ is a collection of graph $\mathcal{K}_{i}$ for disease category $i$. $\mathcal{K}_{i}$ is based on the diagnosis criteria in existing guidelines (refer to supplementary material for details). $\mathcal{K}_{i}$’s nodes are either premise $p\in\mathcal{P}_{i}$ (medical statement, e.g., Headache is a symptom of) and diagnoses $d\in\mathcal{D}_{i}$ (e.g., Suspected Stroke). $\mathcal{K}_{i}$ consists of two different types of edges. One is premise-to-diagnosis edges $\mathcal{S}_{i}=\{(p,d)\}$, where $p\in\mathcal{P}_{i}$ and $d\in\mathcal{D}_{i}$; an edge is from $p$ to $d$. This edge represents the necessary premise $p$ to make a diagnosis $d$. We refer to them as supporting edges. The other is diagnosis-to-diagnosis edges $\mathcal{F}_{i}=\{(d,d^{\prime})\}$, where $d,d^{\prime}\in\mathcal{D}_{i}$ and the edge is from $d$ to $d^{\prime}$, which represents the diagnostic flow. These edges are referred to as procedural edges.
A disease category is defined according to an existing guideline, which starts from a certain diagnosis; therefore, a procedural graph $\mathcal{G}_{i}=(\mathcal{D}_{i},\mathcal{F}_{i})$ has only one root node and arbitrarily branches toward multiple leaf nodes that represent PDDs (i.e., the clinical notes in DiReCT are chosen to cover all leaf nodes of $\mathcal{G}_{i}$). Thus, $\mathcal{G}_{i}$ is a tree. We denote the set of the leaf nodes (or PDDs) as $\mathcal{D}^{\star}_{i}\subset\mathcal{D}_{i}$. The knowledge graph is denoted by $\mathcal{K}_{i}=(\mathcal{D}_{i},\mathcal{P}_{i},\mathcal{S}_{i},\mathcal{F}_{i})$.
Figure 2: A part of $\mathcal{K}_{i}$ for $i$ being Acute Coronary Syndromes.
Figure 2 shows a part of $\mathcal{K}_{i}$, where $i$ is Acute Coronary Syndromes (ACS). Premises in $\mathcal{P}_{i}$ and diagnoses in $\mathcal{D}_{i}$ are given in the blue and gray boxes, while PDDs in $\mathcal{D}^{\star}_{i}$ are ones without outgoing edges (i.e., STEMI-ACS and NSTEMI-ACS, and UA). The black and red arrows are edges in $\mathcal{S}$ and $\mathcal{F}$, respectively, where the black arrows indicate the supporting edges.
@@ 3.3 Data Annotation
Let $d^{\star}\in\mathcal{D}^{\star}_{i}$ denote the PDD of disease category $i$ associated with $R$. We can find a subgraph $\mathcal{K}_{i}(d^{\star})$ of $\mathcal{K}_{i}$ that contains all ancestors of $d^{\star}$, including premises in $\mathcal{P}_{i}$. We also denote the set of supporting edges in $\mathcal{K}_{i}(d^{\star})$ as $\mathcal{S}_{i}(d^{\star})$. Our annotation process is, for each supporting edge $(p,d)\in\mathcal{S}_{i}(d^{\star})$, to extract observation $o\in\mathcal{O}$ in $R$ (highlighted text in the clinical note in Figure 3) and provide rationalization $z$ of this deduction why $o$ is a support for $d$ or corresponds to $p$ [^3] All annotations strictly follow the procedural flow in $\mathcal{K}_{i}$, and each observation is only related to one diagnostic node. If $R$ does not provide sufficient observations for the PDD (which may happen when a certain test is omitted), the annotators were asked to add plausible observations to $R$. This choice compromises the fidelity of our dataset to the original clinical notes, but we chose it for the completeness of the dataset. They form the explanation =$\mathcal{E}=\{(o,z,d)\}$ for $(R,d^{\star})$. This annotation process was carried out by 9 clinical physicians and subsequently verified for accuracy and completeness by three senior medical experts.
@@ 3.5 Evaluation Metrics
(3) Faithfulness of explanations Faith evaluates if the diagnostic flow toward the PDD is fully supported by observations with faithful rationalizations. This involves establishing a one-to-one correspondence between deductions in the prediction and the ground truth. We use the correspondences established for computing $\textit{Obs}^{\text{comp}}$. Let $o\in\mathcal{O}$ and $\hat{o}\in\hat{\mathcal{O}}$ denote corresponding observations. This correspondence is considered successful if $z$ and $\hat{z}$ as well as $d$ and $\hat{d}$ associated with $o$ and $\hat{o}$ matches. Let $m(\mathcal{E},\hat{\mathcal{E}})$ denote the number of successful matches. We use the ratio of $m(\mathcal{E},\hat{\mathcal{E}})$ to $|\mathcal{O}\cap\hat{\mathcal{O}}|$ and $|\mathcal{O}\cup\hat{\mathcal{O}}|$ as evaluation metrics $\textit{Exp}^{\text{com}}$ and $\textit{Exp}^{\text{all}}$, respectively, to see failures come from observations or explanations and diagnosis.
@@ 4 Baseline
Figure 4 shows an overview of our baseline with three LLM-based modules narrowing-down, perception, and reasoning (refer to the supplementary material for more details). The narrowing-down module $U$ takes $R$ as input to make a prediction $\hat{i}$ of the disease category, i.e., $\hat{i}=U(R)$.
Let $d_{t}\in\mathcal{D}_{\hat{i}}$ be the diagnosis that has been reached with $t$ iterations over $\mathcal{K}_{\hat{i}}$, where $t$ corresponds to the depth of node $d_{t}$ and so is less than or equal to the depth of $\mathcal{K}_{i}$. $d_{0}$ is the root node of $\mathcal{K}_{\hat{i}}$. For $d_{0}$, we apply the perception module to extract all observations in $R$ and explanation $\mathcal{E}_{0}$ to support $d_{0}$ as
$\hat{\mathcal{O}},\hat{\mathcal{E}_{0}}=W(d_{0},\mathcal{K}_{\hat{i}}).$ (3)
$\mathcal{K}_{\hat{i}}$ is supplied to facilitate the model to extract all observations for the following reasoning process. [^5] We used only pairs of an observation and a premise. We abuse $\mathcal{K}$ to mean this for notation simplicity.
Diagnosis $d_{t}$ identifies the set $\{d_{n}\}_{n}$ of its children and so the set $\mathcal{P}_{\hat{i}}(\{d_{n}\}_{n})=\{p\in\mathcal{P}_{i}|(p,d_{n})\in\mathcal{S}_{i},d_{n}\in\{d_{n}\}\}$ of premises that support $d_{n}$. Therefore, our reasoning module $V$ iteratively and greedily identifies the next step’s diagnosis (i.e., $d_{t+1}$) from $\{d_{n}\}_{n}$, making a rationalization for each deduction. That is, $V$ verifies whether there exist $o$’s in $\hat{\mathcal{O}}$ that supports one $d_{n}$. If $d_{n}$ is fully supported, $d_{n}$ is identified as $d_{t+1}$ for the $(t+1)$-th iteration, i.e.,
$d_{t+1},\hat{\mathcal{E}}_{t+1}=V(\hat{\mathcal{O}},\{d_{n}\},\mathcal{P}_{\hat{i}}(\{d_{n}\}_{n})).$ (4)
Otherwise, the reasoning module fails. $V$ is repeated until $d_{t^{\prime}}$ in $\mathcal{D}^{\star}_{\hat{i}}$ is found or it fails. In our annotation, each observation contributes to deducing only one $d_{t}$. Therefore, if an observation in ^$\hat{\mathcal{E}}_{t+1}$ is included in the preceding sets of explanations ^$\hat{\mathcal{E}}_{0}$ to ^$\hat{\mathcal{E}}_{t}$, the corresponding explanation in the preceding sets is removed.
@@ 5.1 Experimental Setup
5.2 Results
Table 3: Diagnostic reasoning ability of different LLMs under the proposed baseline method.
@@ 5.2 Results
LLama3 70B 0.752 0.540 0.277±${}_{\pm\text{0.146}}$ 0.537±${}_{\pm\text{0.192}}$ 0.256±${}_{\pm\text{0.142}}$ 0.395±${}_{\pm\text{0.320}}$ 0.112±${}_{\pm\text{0.110}}$
@@
GPT-4 turbo 0.772 0.572 0.446±${}_{\pm\text{0.207}}$ 0.491±${}_{\pm\text{0.180}}$ 0.371±${}_{\pm\text{0.186}}$ 0.475±${}_{\pm\text{0.363}}$ 0.199±${}_{\pm\text{0.181}}$
@@
LLama3 70B 0.735 0.581 0.262±${}_{\pm\text{0.146}}$ 0.501±${}_{\pm\text{0.208}}$ 0.236±${}_{\pm\text{0.131}}$ 0.463±${}_{\pm\text{0.374}}$ 0.125±${}_{\pm\text{0.117}}$
@@
GPT-4 turbo 0.781 0.614 0.431±${}_{\pm\text{0.207}}$ 0.458±${}_{\pm\text{0.187}}$ 0.353±${}_{\pm\text{0.170}}$ 0.633±${}_{\pm\text{0.338}}$ 0.247±${}_{\pm\text{0.201}}$
Table 4: Evaluation of diagnostic reasoning ability of LLMs when no external knowledge is provided.

### v3

@@ Abstract
Large language models (LLMs) have recently showcased remarkable capabilities, spanning a wide range of tasks and applications, including those in the medical domain. Models like GPT-4 excel in medical question answering but may face challenges in the lack of interpretability when handling complex tasks in real clinical settings. We thus introduce the diagnostic reasoning dataset for clinical notes (DiReCT), aiming at evaluating the reasoning ability and interpretability of LLMs compared to human doctors. It contains 511 clinical notes, each meticulously annotated by physicians, detailing the diagnostic reasoning process from observations in a clinical note to the final diagnosis. Additionally, a diagnostic knowledge graph is provided to offer essential knowledge for reasoning, which may not be covered in the training data of existing LLMs. Evaluations of leading LLMs on DiReCT bring out a significant gap between their reasoning ability and that of human doctors, highlighting the critical need for models that can reason effectively in real-world clinical scenarios Data and code are available at https://github.com/wbw520/DiReCT..
@@ 1 Introduction
Despite the advancements, interpretability is critical, particularly in medical NLP tasks [Liévin et al., 2024] because these tasks directly impact patient health and treatment decisions. Without clear interpretability, there’s a risk of misdiagnosis and improper treatment, making it vital for ensuring medical safety. Some studies assess this capability over medical QA [Pal et al., 2022, Li et al., 2023, Chen et al., 2024] or natural language inference (NLI) [Jullien et al., 2023]. Putting more attention on interpretability, they use relatively simple tasks as testbeds, taking short text as input. Nevertheless, real-world clinical tasks are often more complex [Gao et al., 2023a], as illustrated in Figure 1, a typical diagnosis requires comprehending and combining various information, such as health records, physical examinations, and laboratory tests, for further reasoning of possible diseases in a step-by-step manner following the established guidelines. This observation suggests that both perception, or reading (e.g., finding necessary information in the medical record) and reasoning (determining the disease based on the observations) should be counted when evaluating interpretability in LLM-based medical NLP tasks.
@@
To underscore the challenge offered by our dataset, we propose a simple AI-agent based baseline [Xi et al., 2023, Tang et al., 2023] that utilizes the knowledge graph to decompose the diagnosis into a sequence of diagnoses from a smaller number of observations. Our experimental findings indicate that current state-of-the-art LLMs still fall short of aligning well with human doctors.
Contribution. DiReCT offers a new challenge in diagnosis from a complex clinical note with explicit knowledge of established guidelines. This challenge aligns with a realistic medical scenario that doctors are experiencing. In the application aspect, the dataset facilitates the development of a model to support doctors in diagnosis, which is error-prone [Middleton et al., 2013, Liu et al., 2022]. From the technical aspect, the dataset can benchmark models’ ability to read long text and find necessary observations for multi-evidence entailment tree reasoning, an extension of the original entailment tree explanation [Dalvi et al., 2021] for complex scenarios in medical NLP tasks. As shown in Figure 3, this is not trivial because of the variations in writing; superficial matching does not help, and medical knowledge is vital. Meanwhile, reasoning itself is facilitated by the knowledge graph. The model does not necessarily have the knowledge of diagnostic guidelines. With this choice, the knowledge graph explains the reasoning process, which is also beneficial when deploying such a diagnosis assistant system in practical uses.
@@ 3.2 Diagnostic Knowledge Graph
Our knowledge graphs $\mathcal{K}=\{k_{i}\}$ is a collection of graph $k_{i}$ for disease category $i$. $k_{i}$ is based on the diagnosis criteria in existing guidelines (refer to supplementary material for details). $k_{i}$’s nodes are either premise $p\in\mathcal{P}_{i}$ (medical statement, e.g., Headache is a symptom of) and diagnoses $d\in\mathcal{D}_{i}$ (e.g., Suspected Stroke). $k_{i}$ consists of two different types of edges. One is premise-to-diagnosis edges $\mathcal{S}_{i}=\{(p,d)\}$; an edge is from $p$ to $d$. This edge represents the necessary premise $p$ to make a diagnosis $d$. We refer to them as supporting edges. The other is diagnosis-to-diagnosis edges $\mathcal{F}_{i}=\{(d,d^{\prime})\}$, where $d,d^{\prime}\in\mathcal{D}_{i}$ and the edge is from $d$ to $d^{\prime}$, which represents the diagnostic flow. These edges are referred to as procedural edges.
A disease category is defined according to an existing guideline, which starts from a certain diagnosis; therefore, a procedural graph $g_{i}=(\mathcal{D}_{i},\mathcal{F}_{i})$ ($\mathcal{G}=\{g_{i}\}$) has only one root node and arbitrarily branches toward multiple leaf nodes that represent PDDs (i.e., the clinical notes in DiReCT are chosen to cover all leaf nodes of $g_{i}$). Thus, $g_{i}$ is a tree. We denote the set of the leaf nodes (or PDDs) as $\mathcal{D}^{\star}_{i}\subset\mathcal{D}_{i}$. The knowledge graph is denoted by $k_{i}=(\mathcal{D}_{i},\mathcal{P}_{i},\mathcal{S}_{i},\mathcal{F}_{i})$.
Figure 2: A part of $k_{i}$ for $i$ being Acute Coronary Syndromes.
Figure 2 shows a part of $k_{i}$, where $i$ is Acute Coronary Syndromes (ACS). Premises in $\mathcal{P}_{i}$ and diagnoses in $\mathcal{D}_{i}$ are given in the blue and gray boxes, while PDDs in $\mathcal{D}^{\star}_{i}$ are ones without outgoing edges (i.e., STEMI-ACS and NSTEMI-ACS, and UA). The black and red arrows are edges in $\mathcal{S}$ and $\mathcal{F}$, respectively, where the black arrows indicate the supporting edges.
@@ 3.3 Data Annotation
Let $d^{\star}\in\mathcal{D}^{\star}_{i}$ denote the PDD of disease category $i$ associated with $R$. We can find a subgraph $k_{i}(d^{\star})$ of $k_{i}$ that contains all ancestors of $d^{\star}$, including premises in $\mathcal{P}_{i}$. We also denote the set of supporting edges in $k_{i}(d^{\star})$ as $\mathcal{S}_{i}(d^{\star})$. Our annotation process is, for each supporting edge $(p,d)\in\mathcal{S}_{i}(d^{\star})$, to extract observation $o\in\mathcal{O}$ in $R$ (highlighted text in the clinical note in Figure 3) and provide rationalization $z$ of this deduction why $o$ is a support for $d$ or corresponds to $p$ [^3] All annotations strictly follow the procedural flow in $k_{i}$, and each observation is only related to one diagnostic node. If $R$ does not provide sufficient observations for the PDD (which may happen when a certain test is omitted), the annotators were asked to add plausible observations to $R$. Refer to amended data points in supplementary for details. They form the explanation =$\mathcal{E}=\{(o,z,d)\}$ for $(R,d^{\star})$. This annotation process was carried out by 9 clinical physicians and subsequently verified for accuracy and completeness by three senior medical experts.
@@ 3.5 Evaluation Metrics
(3) Faithfulness of explanations evaluates if the diagnostic flow toward the PDD is fully supported by observations with faithful rationalizations. This involves establishing a one-to-one correspondence between deductions in the prediction and the ground truth. We use the correspondences established for computing $\textit{Obs}^{\text{comp}}$. Let $o\in\mathcal{O}$ and $\hat{o}\in\hat{\mathcal{O}}$ denote corresponding observations. This correspondence is considered successful if $z$ and $\hat{z}$ as well as $d$ and $\hat{d}$ associated with $o$ and $\hat{o}$ matches. Let $m(\mathcal{E},\hat{\mathcal{E}})$ denote the number of successful matches. We use the ratio of $m(\mathcal{E},\hat{\mathcal{E}})$ to $|\mathcal{O}\cap\hat{\mathcal{O}}|$ and $|\mathcal{O}\cup\hat{\mathcal{O}}|$ as evaluation metrics $\textit{Exp}^{\text{com}}$ and $\textit{Exp}^{\text{all}}$, respectively, to see failures come from observations or explanations and diagnosis.
@@ 4 Baseline
Figure 4 provides an overview of our baseline, which comprises three LLM-based modules: narrowing-down ($U$), perception ($W$), and reasoning ($V$). In our experiments, each module utilizes the same type of LLM with different prompts (refer to the supplementary material for more details). $U$ analyze the entire note $R$ to determine the possible disease type $\hat{i}$. $W$ extracts observations that may lead to diseases from each $r$, producing a list of original disease descriptions. $V$ iteratively derives possible diseases from observations based on the diagnosis knowledge graph, providing rationales for each deduction $(o,z,d)$.
The narrowing-down module $U$ takes $R$ as input to make a prediction $\hat{i}$ of the disease category, i.e., $\hat{i}=U(R)$. Let $d_{t}\in\mathcal{D}_{\hat{i}}$ be the diagnosis that has been reached with $t$ iterations over $k_{\hat{i}}$, where $t$ corresponds to the depth of node $d_{t}$ and so is less than or equal to the depth of $k_{\hat{i}}$. $d_{0}$ is the root node of $k_{\hat{i}}$. For $d_{0}$, we apply the perception module to extract all observations in $R$ and explanation $\mathcal{E}_{0}$ to support $d_{0}$ as
$\hat{\mathcal{O}},\hat{\mathcal{E}_{0}}=W(d_{0},k_{\hat{i}}).$ (3)
$k_{\hat{i}}$ is supplied to facilitate the model to extract all observations for the following reasoning process. [^5] We used only pairs of an observation and a premise. We abuse $\mathcal{K}$ to mean this for notation simplicity. The perception model can also utilize $g_{i}$ instead of $k_{i}$ for the first task. After the perception module $W$ (iteration $t=0$), we obtain all observations $\mathcal{\hat{O}}$, the root node of the diagnosis $d_{0}$, and an explanation ^$\hat{\mathcal{E}}_{0}$ for the initial iteration. Assuming that by iteration $t$, we already know the diagnosis for iteration $t$ as $d_{t}$. $\{d_{n}\}$ is the set of $d_{t}$’s children, and $\mathcal{P}_{\hat{i}}(\{d_{n}\})$ represents the corresponding premises that support each $d_{n}$. $V$ identifies the diagnosis for the next step, $d_{t+1}$, and provides a justification $\mathcal{E}_{t+1}$. $V$ will verify if there is any $\hat{o}$ in $\mathcal{\hat{O}}$ that supports a $d_{n}$. If fully supported, $d_{n}$ is identified as $d_{t+1}$ for the $(t+1)$-th iteration, i.e.,
$d_{t+1},\hat{\mathcal{E}}_{t+1}=V(\hat{\mathcal{O}},\{d_{n}\},\mathcal{P}_{\hat{i}}(\{d_{n}\})),$ (4)
$V$ continues until $d_{t+1}$ in $\mathcal{D}^{*}$ is identified. If no observation supports a $d_{n}$, the reasoning process will be stopped.
In our annotation, an observation $o$ is associated with only one $d$. However, our method employs an iterative reasoning pipeline. Initially, the perception module $W$ generates an explanation set ^$\hat{\mathcal{E}}_{0}$, linking all $\hat{o}$ to $d_{0}$. During the $t$-th iteration of $V$, the explanation set is ^$\hat{\mathcal{E}}_{t}$, where at least one $\hat{o}$ is linked to $d_{t}$. The final diagnosis explanation is the combination of ^${\hat{\mathcal{E}}_{0},\dots,\hat{\mathcal{E}}_{T}}$ and ${d_{0},\dots,d_{T}}$, where $T$ represents the final iteration. In this combination, if an $\hat{o}$ is eventually processed in the iteration for ^$\hat{\mathcal{E}}_{t}$, the corresponding $(o,z,d)$ in all preceding ^${\hat{\mathcal{E}}_{0},\dots,\hat{\mathcal{E}}_{t-1}}$ will be removed. That is, $\hat{o}$ will always be possessed by the $d_{t}$ closest to the leaf PDD node.
@@ 5.1 Experimental Setup
Table 3: Evaluation of diagnostic reasoning ability using $\mathcal{G}$ or $\mathcal{K}$ as input.
@@
LLama3 70B 0.822 0.606 0.306±${}_{\pm\text{0.151}}$ 0.543±${}_{\pm\text{0.183}}$ 0.279±${}_{\pm\text{0.146}}$ 0.409±${}_{\pm\text{0.328}}$ 0.124±${}_{\pm\text{0.120}}$
@@
GPT-4 turbo 0.804 0.601 0.486±${}_{\pm\text{0.207}}$ 0.481±${}_{\pm\text{0.180}}$ 0.391±${}_{\pm\text{0.189}}$ 0.481±${}_{\pm\text{0.362}}$ 0.210±${}_{\pm\text{0.188}}$
@@
LLama3 70B 0.795 0.651 0.268±${}_{\pm\text{0.147}}$ 0.524±${}_{\pm\text{0.211}}$ 0.258±${}_{\pm\text{0.142}}$ 0.549±${}_{\pm\text{0.372}}$ 0.152±${}_{\pm\text{0.130}}$
@@
GPT-4 turbo 0.808 0.611 0.470±${}_{\pm\text{0.209}}$ 0.459±${}_{\pm\text{0.190}}$ 0.382±${}_{\pm\text{0.192}}$ 0.645±${}_{\pm\text{0.385}}$ 0.273±${}_{\pm\text{0.216}}$
5.2 Results
[…]
Table 4: Evaluation of diagnostic reasoning ability without external knowledge.